The area of a triangle is 12 inches. The height is 1/2 of the base minus 1. What is the base of the triangle?
Answer by venugopalramana(3286)

The area of a triangle is 12 inches. The height is 1/2 of the base minus 1. What is the base of the triangle?
LET BASE =B
HALF OF BASE =B/2
HEIGHT =B/2 -1=(B-2)/2
AREA=(1/2)*B*(B-2)/2=12
B^2-2B=4*12=48
B^2-2B-48=0
B^2-8B+6B-48=0
B(B-8)+6(B-8)=0
(B-8)(B+6)=0
B=8....AND H=8/2 -1=3